# Intelligent Framework for Human-Robot Collaboration: Safety, Dynamic Ergonomics, and Adaptive Decision-Making

### v1

@@ III-H OpenPose: Pose Detection
III-I SlowOnly: Temporal Analysis
The SlowOnly model, a three-dimensional convolutional network (3D ConvNet), analyzes the slow, repetitive movements of operators. The input is a sequence of video frames $F_{text{in}}$, sampled at regular intervals:
@@ III-I SlowOnly: Temporal Analysis
where $I_{t}$ is the video frame at time $t$, $\Delta t$ the temporal sampling interval, and $T$ the total number of frames in the sequence.
Three-dimensional convolution applied to this sequence extracts the temporal and spatial features:
$F_{\text{out}}=\text{Conv3D}(F_{\text{in}},K,T)$
III-J OWAS-Based Ergonomic Assessment
The Ovako Working Posture Analysis System (OWAS), classifies operators’ postures into risk categories, allowing the system to monitor and intervene when postures are detected that could lead to musculoskeletal injuries.
III-K Classification of Postures
The classification of operator postures, in Fig.3, involves four categories:
@@ III-K Classification of Postures
    Posture of the back: straight, flexed or rotated.
@@
    Posture of the arms: above or below shoulder level.
@@
    Position of legs: standing, sitting, kneeling or squatting.
@@
    Weight lifted: amount of load lifted by the operator.
Each combination of these categories is ranked according to an ergonomic risk score, which indicates the level of urgency for corrective action:
  • Level 1: No ergonomic risk, no need for intervention.
  • Level 2: Moderate risk, long-term corrective interventions may be needed.
  • Level 3: High risk, short-term corrective intervention required.
  • Level 4: Extremely risky, immediate corrective action is required.
III-L Calculation of OWAS Score
The overall OWAS score for a posture is calculated as a weighted combination of the scores of the different categories (back, arms, legs, and load). The formula for calculating the overall score is:
@@ III-L Calculation of OWAS Score
  • $P_{\text{back}}$, $P_{\text{arm}}$, $P_{\text{leg}}$, $P_{\text{weight}}$ are the scores assigned to each category,
  • $\alpha$, $\beta$, $\gamma$, $\delta$ are the weights assigned according to the severity of risk for each category.
Figure 4: Behaviour Tree diagram for dynamic role allocation between robot and operator based on physical and operational conditions..
III-M Integration of OWAS into the Framework
OWAS is integrated directly into the framework to monitor operators’ posture in real time, using data from OpenPose and SlowOnly. Data from body joints, collected by OpenPose, are mapped to OWAS categories (back, arms, legs, and load) to automatically assess operator posture.
The assessment process consists of the following steps:
@@ III-M Integration of OWAS into the Framework
    Postural data collection: OpenPose detects the position of the body joints and builds a virtual skeleton.
@@
    Posture Classification: The collected information is mapped to OWAS categories.
@@
    Calculation of OWAS score: The overall OWAS score is calculated for each combination of detected postures and loads.
@@
    Dynamic Intervention: If the OWAS score exceeds a predefined threshold, the system suggests the operator to change the posture or stop the activity to prevent injury.
III-N Behaviour Tree (BT)
based automation. Within the proposed framework, dynamic role allocation between human operator and robot is managed through a Behaviour Tree (BT). This decision structure allows the system to adapt robotic behavior based on operational conditions, operator actions, and real-time sensory inputs. Unlike traditional approaches, such as finite-state diagrams or Markovian decision models, Behaviour Tree offers greater flexibility and modularity, which is particularly relevant in dynamic industrial settings where human-robot collaboration requires fast and accurate decisions.
III-O Behaviour Tree structure
The Behaviour Tree implemented in the framework is hierarchically structured and composed of different types of nodes:
  • Action Nodes: They perform concrete operations, such as “grabbing an object” or “moving to a destination.” These nodes allow the robot to perform the required actions based on the detected conditions.
@@ III-O Behaviour Tree structure
    • Sequence (Sequence): Nodes are executed in sequential order until all actions are completed correctly. If a node fails, execution stops.
@@
Figure 4 illustrates the structure of the decision tree used in the framework to dynamically allocate roles between the operator and the robot based on operational conditions and the ergonomic situation.
This modular structure allows the system to perform complex actions and make dynamic decisions based on conditions sensed in real time. For example, if the human operator is fatigued or has not completed a certain sequence of movements, the BT automatically triggers alternatives such as robot intervention to lift and move the package.In this way, the BT manages role allocation flexibly, maintaining operator safety and operational efficiency.
III-P Execution of the Behaviour Tree
The execution of the Behaviour Tree starts from the root node, which manages the operational flow of the robot. The root sequence starts a series of conditional and action nodes, including:
@@ III-P Execution of the Behaviour Tree
  • GoHoming: Verifies that the robot is in the correct starting position before starting any action.
@@
  • CheckHumanSequenceCompleted: Evaluates whether the operator has completed the intended sequence of movements.
@@
  • CheckBiomechanicsFatigue: Monitors operator biomechanics fatigue based on ergonomic data collected in real time.
If all conditions are met, the BT proceeds with a series of specific actions, such as:
  • MoveToParcel: The robot moves toward the parcel to be lifted.
  • GraspParcel: It performs the action of grabbing the parcel.
  • MoveToContainer: Moves the parcel to the designated container.
  • ReleaseParcel: Releases the parcel into the container.
In case the ergonomic condition of the operator is not optimal, such as due to muscle fatigue or poor posture, the BT triggers corrective actions, such as suggesting the operator to stop the activity or take a secondary role. This ensures safe and efficient interaction between human and robot.
III-Q Advantages of using Behaviour Tree in the framework
@@
The adoption of the Behaviour Tree in the implemented framework offers several advantages over other control techniques, including:
@@
  • Modularity: Each node in the BT is independent and modifiable, allowing behaviors to be added or removed without affecting the entire system.
  • Flexibility: The system can dynamically adapt to changes in the operator’s physical condition or variations in operational activities, responding in real time to factors such as musculoskeletal fatigue or variation in load weight.
@@
  • Real-Time Responsiveness: Through integration with visual perception modules (e.g., YOLO and OpenPose) and ergonomic evaluation, the BT enables the robot to make quick and accurate decisions, optimizing both operator safety and operational efficiency.
@@ IV Experiments and Results
IV-A Experiment: Grasping Action Detection and Volumetric Weight Estimation with Variable Size Parcels
TABLE II: Global performance in terms of mAP
. Model mAP@50 (%) mAP@50:95 (%) YOLOv8x-seg 77.6 71.2 YOLOv9-seg 77.7 72.4 YOLO11x-seg 77.8 72.4
TABLE III: Performance per class: accuracy, recall, and mAP
Class Metrics YOLOv8x-seg (%) YOLOv9-seg (%) YOLO11x-seg (%)
Parcel Precision 88.0 85.6 89.8
Recall 79.3 80.9 81.1
mAP@50 89.4 89.4 89.3
mAP@50:95 86.7 86.0 86.8
Hand Precision 89.8 89.6 89.8
Recall 59.9 60.9 61.1
mAP@50 65.7 66.1 66.3
mAP@50:95 57.3 53.2 58.0
Three segmentation models were analyzed: YOLOv8x-seg, YOLOv9-seg e YOLO11x-seg, trained with the same configuration parameters and evaluated on two main classes, , ”parcel” and ”hand”. As shown in Table II, YOLOv8x-segachieved a mAP@50 of 77.6% and a mAP@50:95 of 71.2%, demonstrating a good balance between precision and recall. YOLOv9-seg slightly improved the mAP@50 (77.7%) and maintained the mAP@50:95 at 72.4%, with more efficient latency management. However, YOLO11x-seg outperformed both, achieving a mAP@50 of 77.8% and a mAP@50:95 of 72.4%, showing greater robustness in complex scenarios.
For grasping intention recognition, 40 trials were conducted on dynamic sequences, combining the segmentation model with the probabilistic Kalman Unscented (UKF) filter. The system correctly identified grasping in 37 cases out of 40, with an overall accuracy of 92.5%. For volumetric estimation, 4 parcels of varying sizes and shapes were tested, each in 10 trials for a total of 40 measurements. The estimated volumetric weights, obtained from the segmented masks and depth data provided by the RealSense D435 sensor, were compared with the actual weights. Percentage errors were calculated using the formula:
$E_{p}=\frac{|P_{v,\text{st}}-P_{v,\text{re}}|}{P_{v,\text{re}}}\times 100$
where $P_{v,\text{st}}$ e $P_{v,\text{re}}$ represent estimated and actual volumetric weights, respectively. The overall mean percent error over all trials was calculated as 17.58%, with greater variability observed in the extreme size parcels, attributable to the limitations of the sensor in terms of spatial resolution and surface reflectivity.
In conclusion, the model comparison confirmed that YOLO11x-seg represents the best performing solution for the operational context considered. Its combination of robustness, accuracy, and ability to handle difficult conditions, such as detection of moving or partially occluded hands, makes it ideal for grasping and volumetric estimation applications. These results highlight the advantages of integrating a deep learning model with probabilistic tracking algorithms, demonstrating superior robustness compared to the isolated use of neural networks. However, volumetric weight estimation showed sensitivity to the quality of depth data provided by the RealSense D435 sensor, with a reduction in accuracy due to limitations in spatial resolution or surface reflectivity. This suggests that further improvements could be achieved through refinement of sensor calibration or integration of data from multiple sensors to compensate for current limitations.
IV-B Experiment: Ergonomic Monitoring, Risk Classification and Robotic Intervention via Behavior Tree
The objective of this experiment is to evaluate the framework’s ability to monitor operator postures in real time, classify ergonomic risk, and manage task transfer to the robot in the presence of critical conditions. During testing, the operator performed movements typical of an industrial environment, such as bending and lifting, while the system monitored posture using OpenPose and classified ergonomic risk using the Ovako Working Posture Analysis System (OWAS) method.
When medium-high-risk (OWAS class 3) or high-risk (OWAS class 4) postures were detected, or volumetric weight in excess of allowable limits, the system generated a biomechanical fatigue message. This message was processed by a Behavior Tree (BT), which combined ergonomic data, action recognition results, and volumetric weight values to determine whether robotic intervention should be activated. This approach ensured modular and responsive decision making, allowing a smooth transition between the operator and the robot.
The data collected during the experiment showed that 12,8% of the postures belonged to OWAS class 1 (no risk), l’84,6% belonged to OWAS class 2 (moderate risk), no postures were classified in OWAS class 3 (medium-high risk), and 2,6% belonged to OWAS class 4 (high risk), as shown in Fig. 6. These results confirm the system’s ability to distinguish between ergonomically safe and hazardous postures, demonstrating the effectiveness of the framework in monitoring and classifying risk.
Temporally, the system maintained an average update rate of 14.99 Hz, with an average latency of 0.57 seconds for pose monitoring, as shown in Fig. 6. These values ensured that any postural changes were detected and analyzed in a timely manner. In addition, the average response time to risky situations was 0.07 seconds, as shown in Fig. 8, demonstrating the efficiency of the framework in identifying critical conditions and responding quickly.
These results highlight not only the framework’s efficiency in detecting and analyzing critical conditions, but also its role in supporting a seamless operational transition. The reduced overall latency, as calculated through pose monitoring and response time to risky situations, ensured that the transfer of activities to the robot was smooth and responsive. This level of operational readiness, combined with the system’s ability to integrate ergonomic and actionable data, represents a significant step toward optimizing human-robot collaboration.

### v2

@@ III-B1 Pose Detection with OpenPose
III-C Action Recognition Module
The Action Recognition Module analyzes the operator’s movements over time to identify specific actions and patterns, contributing to both ergonomic assessment and decision-making.
III-C1 Temporal Analysis with SlowOnly
The SlowOnly model, a specialized three-dimensional convolutional network (3D ConvNet), analyzes the slow, repetitive movements of operators. The input is a sequence of video frames $F_{\text{in}}$, sampled at regular temporal intervals:
@@ III-C1 Temporal Analysis with SlowOnly
Where:
  • $I_{t}\in\mathbb{R}^{H\times W\times C}$ represents the video frame at time $t$, with $H$ being the height, $W$ the width, and $C$ the number of channels of the frame
  • $\Delta t\in\mathbb{R}^{+}$ is the temporal sampling interval between consecutive frames
  • $T\in\mathbb{N}$ is the total number of frames in the sequence
Three-dimensional convolution applied to this sequence extracts both spatial and temporal features:
$F_{\text{out}}=\text{Conv3D}(F_{\text{in}},K,S)$
Where:
  • $F_{\text{out}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D\times N}$ is the output feature tensor
  • $K\in\mathbb{R}^{k_{h}\times k_{w}\times k_{t}\times C\times N}$ represents the 3D convolutional kernels
  • $S\in\mathbb{N}^{3}$ denotes the stride parameters $(s_{h},s_{w},s_{t})$
The temporal features extracted by SlowOnly are then forwarded to the Ergonomic Assessment Module for evaluation of the operator’s movements using the OWAS methodology.
III-D Ergonomic Assessment Module
The Ergonomic Assessment Module integrates data from the previous modules to evaluate the operator’s posture and working conditions according to established ergonomic standards.
III-D1 Ovako Working Posture Analysis System
The Ovako Working Posture Analysis System (OWAS) classifies operators’ postures into risk categories, allowing the system to monitor and intervene when postures are detected that could lead to musculoskeletal injuries.
Fig. 3 illustrates the OWAS evaluation matrix used in our framework. This standardized tool provides a systematic approach to assessing ergonomic risks through a color-coded matrix that indicates different levels of necessary intervention based on various posture combinations.
Classification of Postures
The classification of operator postures, as shown in Fig. 3, is based on a systematic coding system that evaluates four key body segments:
@@ III-D1 Ovako Working Posture Analysis System
    Posture of the back (1-4):
    • (1) Straight
    • (2) Bent
    • (3) Twisted
    • (4) Bent and twisted
@@
    Posture of the arms (1-3):
    • (1) Both arms below shoulder level
    • (2) One arm at or above shoulder level
    • (3) Both arms at or above shoulder level
@@
    Position of legs (1-7):
    • (1) Sitting
    • (2) Standing on two straight legs
    • (3) Standing on one straight leg
    • (4) Standing or squatting on two bent legs
    • (5) Standing or squatting on one bent leg
    • (6) Kneeling
    • (7) Walking
@@
    Weight lifted (1-3):
    • (1) Less than or equal to 10 kg
    • (2) Greater than 10 kg and less than or equal to 20 kg
    • (3) Greater than 20 kg
The OWAS evaluation process combines these four codes to determine a risk level using the standardized evaluation matrix shown in Fig. 3. This matrix cross-references specific combinations of postures and load to assign one of four action categories:
  • Action Category 1: No action required - normal posture without harmful effect on the musculoskeletal system.
  • Action Category 2: Corrective actions required in the near future - posture has some harmful effect on the musculoskeletal system.
  • Action Category 3: Corrective actions should be done as soon as possible - posture has a distinctly harmful effect on the musculoskeletal system.
  • Action Category 4: Corrective actions for improvement required immediately - posture has an extremely harmful effect on the musculoskeletal system.
Figure 4: Hierarchical structure of the implemented Behaviour Tree for dynamic role allocation. The tree starts with a ROOT Fallback node and branches through a Main Sequence into conditional checks for robot position, human sequence completion, biomechanical fatigue, and parcel weight. Based on these evaluations, the Action Sequence Fallback node directs execution either to robot-led actions (via No Walk With Parcel Sequence) or human-led operations with robot support (via Walk With Parcel Sequence).
OWAS Score Calculation
While the standard OWAS methodology determines the action category directly through the evaluation matrix, our implementation calculates a continuous weighted score to enable more nuanced ergonomic assessment. The formula for calculating this weighted OWAS score is:
@@
  • $P_{\text{back}}$, $P_{\text{arm}}$, $P_{\text{leg}}$, $P_{\text{weight}}$ are the numerical codes assigned to each postural category (ranging from 1-4, 1-3, 1-7, and 1-3 respectively)
  • $\alpha$, $\beta$, $\gamma$, $\delta$ are weighting coefficients that modulate the relative contribution of each postural factor
The weighting coefficients ($\alpha$, $\beta$, $\gamma$, $\delta$) are derived from ergonomic research on work-related musculoskeletal disorders. In the original OWAS methodology [34], these weights reflect the relative impact of different body segments on overall postural strain. The exact values are calibrated based on:
  • The prevalence of disorders in different body regions in industrial settings
  • The biomechanical load associated with each posture type
  • The time sustainability of various postures before discomfort onset
  • The recovery time needed after maintaining specific postures
The calculated weighted score is then mapped to the four action categories using threshold values that align with the OWAS evaluation matrix shown in Fig. 3. This approach allows our system to continuously monitor ergonomic risk levels and trigger appropriate interventions when hazardous postures are detected.
Framework Integration of Ergonomic Assessment
The OWAS methodology is integrated directly into the framework to monitor operators’ posture in real time, establishing a computational pathway between sensor data and ergonomic risk assessment. The integration process consists of the following steps:
@@
    Skeletal data acquisition: OpenPose detects the 3D coordinates and confidence scores of key body joints, while SlowOnly analyzes temporal movement patterns to identify sustained postures.
@@
    Posture mapping: The acquired skeletal data is mapped to the corresponding OWAS coding system (back, arms, legs), while load parameters are derived from the volumetric calculations performed by the Visual Perception Module.
@@
    OWAS score computation: The weighted score is calculated using the formula previously described, with the calibrated coefficients reflecting the biomechanical impact of each postural component.
@@
    Dynamic intervention: When the computed ergonomic risk exceeds the predefined thresholds corresponding to OWAS action categories, the system initiates appropriate intervention strategies to mitigate the identified hazards.
This approach transforms the traditional manual OWAS assessment into an automated, continuous monitoring system capable of detecting ergonomic hazards in real-time industrial scenarios.
III-E Decision-Making with Behaviour Trees
Within the proposed framework, dynamic role allocation between human operator and robot is managed through a BT. This decision structure allows the system to adapt robotic behaviour based on operational conditions, operator actions, and real-time sensory inputs. Unlike traditional approaches, such as finite-state diagrams or Markovian decision models, the BT offers greater flexibility and modularity, which is particularly relevant in dynamic industrial settings where human-robot collaboration requires fast and accurate decisions.
III-E1 Structure and Execution of the Behaviour Tree
The BT implemented in the framework is hierarchically structured and composed of different types of nodes:
  • Action Nodes: They perform concrete operations, such as ”grabbing an object” or ”moving to a destination.” These nodes allow the robot to perform the required actions based on the detected conditions.
@@ III-E1 Structure and Execution of the Behaviour Tree
    • Sequence: Nodes are executed in sequential order until all actions are completed correctly. If a node fails, execution stops.
@@
As illustrated in Fig. 4, the BT structure begins with a ROOT Fallback node at the top level, which provides the initial decision-making capability. This root node connects to a Main Sequence node that orchestrates the operational flow. From the Main Sequence, the tree branches into five parallel nodes that constitute the core decision-making components:
@@
  • Go Homing (Action): Verifies that the robot is in the correct starting position before initiating operations.
@@
  • Check Human Sequence Completed (Condition): Evaluates whether the operator has completed their intended sequence of movements.
@@
  • Check Biomechanics Fatigue (Condition): Monitors operator biomechanical fatigue based on ergonomic data collected in real time.
  • Check Parcel Weight (Condition): Evaluates whether the weight of the package falls within acceptable parameters for handling.
  • Action Sequence Fallback: A composite fallback node that determines the appropriate action path based on the outcome of preceding conditions.
The execution process begins at the ROOT Fallback node and proceeds through the Main Sequence. Based on the evaluation of the conditional nodes, the tree implements two primary execution pathways:
  • The first pathway through the No Walk With Parcel Sequence directs robot-led actions when the operator’s ergonomic condition indicates risk. This sequence leads to actions including Move To Parcel, Move Arm to Grasp, Correct Grip, Grasp Parcel, Lift Parcel, Move to Container, and Release Parcel.
  • The second pathway through the Walk With Parcel Sequence manages human-led operations with robot support when ergonomic conditions are favorable. This includes monitoring through Execute Human Actions and providing assistance as needed.
When the OWAS assessment indicates elevated ergonomic risk (for example, due to improper posture or fatigue detected via the Check Biomechanics Fatigue node), the BT prioritizes robot intervention. Conversely, if the ergonomic assessment indicates favorable conditions for human operation, the tree allows for human-led execution with robot monitoring. This dynamic allocation ensures that workload is distributed optimally based on real-time ergonomic conditions.
III-E2 Advantages of the Behaviour Tree Approach
@@
The adoption of the BT in the implemented framework offers several advantages over other control techniques, including:
@@
  • Modularity: Each node in the BT is independent and modifiable, allowing behaviours to be added or removed without affecting the entire system. This enables the framework to be easily extended with new actions or conditions as requirements evolve.
  • Flexibility: The system can dynamically adapt to changes in the operator’s physical condition or variations in operational activities. For example, if the human operator shows signs of fatigue or adopts risky postures (as detected by the OWAS assessment), the BT automatically triggers robot intervention to lift and move packages, maintaining operator safety.
@@
  • Real-Time Responsiveness: Through integration with visual perception modules (e.g., YOLO11 and OpenPose) and ergonomic evaluation, the BT enables the robot to make quick and accurate decisions, optimizing both operator safety and operational efficiency. The millisecond-level response capability is essential in dynamic industrial environments where conditions can change rapidly.
TABLE II: Comparative analysis of object detection performance across YOLO model generations
Model Architecture mAP@50 (%) mAP@50:95 (%)
YOLOv8x-seg 77.6 71.2
YOLOv9-seg 77.7 72.4
YOLO11x-seg 77.8 72.4
Note: Mean Average Precision metrics at IoU threshold of 0.5 (mAP@50) and across multiple IoU thresholds from 0.5 to 0.95 (mAP@50:95) demonstrate the incremental improvement trajectory across model generations. Best results are highlighted in bold.
@@ IV Experiments and Results
IV-A Experimental Setup
All experiments were conducted using a standardized hardware and software infrastructure to ensure consistency across different test scenarios:
IV-A1 Hardware Components
  1. 1.
    Visual Sensing: Two Intel RealSense D435 depth cameras were deployed for object detection/tracking and human pose estimation respectively. An additional Logitech C270 webcam was used for capturing video streams for action recognition analysis.
  2. 2.
    Robotic System: The experiments utilized a collaborative robotic arm (Universal Robots UR16) mounted on the Summit XL mobile platform, configured to respond to commands from the decision-making system.
  3. 3.
    Computing Resources: The framework operated on a distributed network comprising four workstations in a ROS-based master-slave configuration. Three workstations were dedicated to vision processing, equipped with different GPUs (NVIDIA GTX 970, NVIDIA GTX 1080Ti, and NVIDIA RTX A3000) to handle various perception tasks. The fourth workstation was the onboard computer of the mobile platform, which managed robot control operations.
IV-A2 Software Framework
The system ran on Ubuntu 20.04 with ROS Noetic providing the communication infrastructure between components. The YOLO11 detection module, OpenPose tracking, and SlowOnly action recognition algorithms were deployed across the vision processing nodes according to their computational requirements. Neural network models were initially trained offline, while experiment execution utilized the distributed GPU setup for real-time inference.
This heterogeneous computing architecture was designed to maintain system responsiveness while processing multiple data streams, enabling the ergonomic assessment and decision-making modules to operate effectively in dynamic conditions.
TABLE III: Class-specific performance evaluation of state-of-the-art YOLO segmentation models
Class Metrics YOLOv8x-seg YOLOv9-seg YOLO11x-seg
Parcel Precision (%) 88.0 85.6 89.8
Recall (%) 79.3 80.9 81.1
mAP@50 (%) 89.4 89.4 89.3
mAP@50:95 (%) 86.7 86.0 86.8
Hand Precision (%) 89.8 89.6 89.8
Recall (%) 59.9 60.9 61.1
mAP@50 (%) 65.7 66.1 66.3
mAP@50:95 (%) 57.3 53.2 58.0
Note: Performance metrics illustrate the differential detection capabilities across object classes, with superior performance observed in parcel detection compared to hand detection. YOLO11x-seg demonstrates improvements in critical metrics, particularly in hand detection precision and recall. Best results for each metric are highlighted in bold.
@@ IV-A2 Software Framework
IV-B Experiment: Grasping Action Detection and Volumetric Weight Estimation with Variable Size Parcels